Muddy Duck Manufacturing (MDM) Inc.’s costing system has two direct cost categories: direct materials and direct manufacturing labour. Manufacturing overhead, both fixed and variable, is allocated to products on the basis of standard direct manufacturing labour hours (DMLH). At the beginning of 2017, MDM adopted the following standards for its manufacturing costs:

Input

Cost per output unit

Direct materials

3kg @ $5 per kg

$15

Direct manufacturing labour

5 hours @ $15 per hour

75

Manufacturing overhead:

Variable

$6 per DMLH

30

Fixed

$8 per DMLH

40

Standard manufacturing cost per output unit

$160

The denominator level for total manufacturing overhead per month in 2017 is 40,000 DMLH. MDM’s flexible budget for February 2017 was based on this denominator level. The records for February indicate the following:

Direct materials purchased = 25,000 kg at $5.20/kg

Direct materials used = 23,100 kg

Direct manufacturing labour = 40,100 hours at $14.60/hour

Total actual manufacturing overhead (variable and fixed) = $600,000

Actual production = 7,800 output units

Required:

1) Prepare a schedule of total standard manufacturing costs for the 7,800 output units in February 2017. (12 marks)

2) For February 2017, calculate the following variances, indicating whether each is favourable (F) or unfavourable (U): (28 marks total with 4 marks each)

a. Direct materials rate variance based on purchases.

b. Direct materials efficiency variance.

c. Direct manufacturing labour rate variance.

d. Direct manufacturing labour efficiency variance.

e. Total manufacturing overhead rate variance.

f. Variable manufacturing overhead efficiency variance.

g. Production-volume variance
Answer #1

1) SCHEDULE OF TOTAL STANDARD MANUFACTURING COST FOR 7800 UNITS

PARTICULAR AMOUNT
DIRECT MATERIAL 7800*3KG*5$ 117000
DIRECT LEBOUR 7800*5H*15$ 585000
MANUFACTURING OVERHEAD
VARIABLE 7800*5*6 DMLH 234000
FIXED 40000*8 320000
TOTALCOST 1256000

2) CALCULATE THE VARIANCE

A) DIRECT MATERIAL RATE VARIANCE

=(23100*5)-(23100*5.20)

= - 4620 (UNFAVOURABLE)

B) DIRECT MATERIAL EFFICIENCY VARIANCE

=(23400*5)-(23100*5)

=1500(FAVOURABLE)

C) DIRECT LABOUR RATE VARIANCE

=(40100*15)-(40100*14.6)

=16040 (FAVOURABLE)

D) DIRECT LABOUR EFFICIENCY VARIANCE

=(39000*15)-(40100*15)

= - 1100(UNFAVOURABE)

E)TOTAL MANUFACTURING OVERHEAD RATE VARIANCE

=554000-600000

= - 46000(UNFAVOURABLE)

F) VARIABLE MANUFACTURING OVERHEAD EFFICIENCY VARIANCE

=6*(40000-39000)

=6000 FAVOURABLE

H) PRODUCTION VOLUME VARIANCE

=8000-7800

=200